((3x-3)/4)+((x+54)/7)=15
Answer by oberobic(2304)

(3x-3)/4 + (x+54)/7 = 15
.
( 7(3x-3) + 4(x+54) ) / 28 = 15
.
7(3x-3) + 4(x+54) = 28*15
.
21x -21 +4x +216 = 420
.
collect and rearrange
.
25x + 195 = 420
.
25x = 420 -195
.
25x = 225
.
x = 225/25
.
x = 9
.
Done.